# Low solution rank of the matrix LASSO under RIP with consequences for rank-constrained algorithms

### v1

@@ 4.3 The idealized nuclear norm subgradient
The solution to this problem is well-known: If $M_{*}$ has SVD $M_{*}=U_{*}\Sigma_{*}V_{*}^{T}$, then the solution to (11) is the singular value soft-thresholded matrix $M_{\lambda}\coloneqq U_{*}\Sigma_{\lambda}V_{*}^{T}$, where, if the singular values of $M_{*}$ are $\sigma_{1},\dots,\sigma_{r^{*}}$,
@@ 4.5 Final rank bounds
Note that
$\min\{r^{*},K\}$ $\leq\max_{\begin{subarray}{c}M\in\mathbf{R}^{d_{1}\times d_{2}}\\ \lVert M\rVert_{\mathrm{F}}=1\\ \operatorname{rank}(M)\leq r^{*}\end{subarray}}~{}\langle\widehat{E}-E_{\lambda},M\rangle^{2}.$
This is true because the right-hand side is the maximum squared Frobenius norm of any rank-$r^{*}$ projection of $\widehat{E}-E_{\lambda}$.
Consider any fixed $M\in\mathbf{R}^{d_{1}\times d_{2}}$ such that $\lVert M\rVert_{\mathrm{F}}=1$ and $\operatorname{rank}(M)\leq r^{*}$. By (13),
$\lvert\langle\widehat{E}-E_{\lambda},M\rangle\rvert\leq\frac{1}{\lambda}\lvert\langle\mathcal{A}^{*}\mathcal{A}(M_{\lambda}-\widehat{M}),M\rangle\rvert+\lvert\langle\mathcal{A}^{*}\mathcal{A}(E_{\lambda})-E_{\lambda},M\rangle\rvert+\frac{1}{\lambda}\lvert\langle\mathcal{A}^{*}(\xi),M\rangle\rvert.$
@@
    $\lvert\langle\mathcal{A}^{*}(\xi),M\rangle\rvert\leq\lVert\mathcal{A}^{*}(\xi)\rVert_{\ell_{2}}\lVert M\rVert_{*}\leq\sqrt{r^{*}}\lVert\mathcal{A}^{*}(\xi)\rVert_{\ell_{2}}.$
@@
    $\lvert\langle\mathcal{A}^{*}\mathcal{A}(E_{\lambda})-E_{\lambda},M\rangle\rvert$ $=\lvert\langle\mathcal{A}(E_{\lambda}),\mathcal{A}(M)\rangle-\langle E_{\lambda},M\rangle\rvert$
    $\leq\delta^{*}\lVert E_{\lambda}\rVert_{\mathrm{F}}\lVert M\rVert_{\mathrm{F}}$
@@
    $\lvert\langle\mathcal{A}^{*}\mathcal{A}(M_{\lambda}-\widehat{M}),M\rangle\rvert$ $\leq\lVert\mathcal{A}(M_{\lambda}-\widehat{M})\rVert\lVert\mathcal{A}(M)\rVert$
@@ 4.6 Uniqueness
(the second inequality is a standard property of RIP constants—see, e.g., [52, Proposition 6.6] for the analogous result in the sparse vector case).
Let $\widehat{M},\widehat{M}^{\prime}$ be optimal solutions to (2). We have shown that $\widehat{M},\widehat{M}^{\prime}$ both have rank at most $r$. Set, as before,
@@
5 Convergence of constrained proximal gradient descent
Theorem 2 is an immediate consequence of the more general result Lemma 3 below, which may be of independent interest (though it and its proof are straightforward modifications of existing results).
Let $\mathcal{C}\subseteq\mathbf{R}^{d}$ be a closed (but not necessarily convex) set. Consider the constrained composite minimization problem
@@ 5 Convergence of constrained proximal gradient descent
If the minimum is not unique, $\mathcal{P}_{\eta}$ can choose one arbitrarily.
@@
In the remainder of this section, we prove Lemma 3. The proof is a modification and generalization of that of [7, Theorem 3], with many ideas (and much of our notation) taken from [53].

### v2

@@ 4.3 The idealized nuclear norm subgradient
As already discussed below Theorem 1, the solution to this problem is well-known: if $M_{*}$ has SVD $M_{*}=U_{*}\Sigma_{*}V_{*}^{T}$, then the unique solution to (11) is the singular value soft-thresholded matrix $M_{\lambda}\coloneqq U_{*}\Sigma_{\lambda}V_{*}^{T}$, where, if the singular values of $M_{*}$ are $\sigma_{1},\dots,\sigma_{r^{*}}$,
@@ 4.5 Final rank bounds
With the convention that singular values are arranged in nonincreasing order, note that
$K\leq\sum_{\ell=1}^{K}\sigma_{\ell}^{2}(\widehat{E}-E_{\lambda}).$
However, we have a circular dependence on $K$. To get something useful, we consider two cases: first, if $K\geq r^{*}$, then $\sigma_{\ell}(\widehat{E}-E_{\lambda})\geq 1$ for all $\ell=1,\dots,r^{*}$, so we certainly have
$r^{*}\leq\sum_{\ell=1}^{r^{*}}\sigma_{\ell}^{2}(\widehat{E}-E_{\lambda}).$
On the other hand, if $K<r^{*}$, we have
$K\leq\sum_{\ell=1}^{K}\sigma_{\ell}^{2}(\widehat{E}-E_{\lambda})\leq\sum_{\ell=1}^{r^{*}}\sigma_{\ell}^{2}(\widehat{E}-E_{\lambda}).$
We can combine these two cases into the inequality
$\min\{r^{*},K\}\leq\sum_{\ell=1}^{r^{*}}\sigma_{\ell}^{2}(\widehat{E}-E_{\lambda}).$ (20)
If we can show that the right-hand side is less than $r^{*}$, we will have a bound on $K$.
Now, to bound the right-hand side of (20), note that the function
$\mathbf{R}^{d_{1}\times d_{2}}\ni A\mapsto\sqrt{\sum_{\ell=1}^{r^{*}}\sigma_{\ell}^{2}(A)}\eqqcolon\lVert A\rVert_{(r^{*})}$
is a matrix norm (in fact, it is a unitarily invariant matrix norm); see, for example, [54, Section 7.4.7] or [55, Section IV.2] for more details on this type of matrix norm.
The important property of $\lVert\cdot\rVert_{(r^{*})}$ for our purposes is that it has the following variational form:
$\lVert A\rVert_{(r^{*})}=\max_{\begin{subarray}{c}B\in\mathbf{R}^{d_{1}\times d_{2}}\\ \lVert B\rVert_{\mathrm{F}}\leq 1\\ \operatorname{rank}(B)\leq r^{*}\end{subarray}}~{}\langle A,B\rangle.$
Indeed, for $B$ of rank at most $r^{*}$, $A^{T}B$ has rank at most $r^{*}$, so
$\langle A,B\rangle$ $=\operatorname{tr}(A^{T}B)$
$\leq\sum_{\ell=1}^{r^{*}}\sigma_{\ell}(A^{T}B)$
$\leq\sum_{\ell=1}^{r^{*}}\sigma_{\ell}(A)\sigma_{\ell}(B)$
$\leq\lVert A\rVert_{(r^{*})}\lVert B\rVert_{\mathrm{F}},$
where the middle inequality is a standard singular value inequality for matrix products, and the last inequality is Cauchy-Schwartz on the singular values. Equality is achieved for $B=\frac{1}{\lVert A\rVert_{(r^{*})}}A_{(r^{*})}$, where $A_{(r^{*})}$ is any (not necessarily unique) rank-$r^{*}$ projection of $A$.
Returning to our problem, we obtain, from (20) and the above considerations,
$\min\{r^{*},K\}\leq\max_{\begin{subarray}{c}B\in\mathbf{R}^{d_{1}\times d_{2}}\\ \lVert B\rVert_{\mathrm{F}}\leq 1\\ \operatorname{rank}(B)\leq r^{*}\end{subarray}}~{}\langle\widehat{E}-E_{\lambda},B\rangle^{2}.$
To bound this, note that for any $B\in\mathbf{R}^{d_{1}\times d_{2}}$ with $\lVert B\rVert_{\mathrm{F}}\leq 1$ and $\operatorname{rank}(B)\leq r^{*}$, we have, by (13),
$\lvert\langle\widehat{E}-E_{\lambda},B\rangle\rvert\leq\frac{1}{\lambda}\lvert\langle\mathcal{A}^{*}\mathcal{A}(M_{\lambda}-\widehat{M}),B\rangle\rvert+\lvert\langle\mathcal{A}^{*}\mathcal{A}(E_{\lambda})-E_{\lambda},B\rangle\rvert+\frac{1}{\lambda}\lvert\langle\mathcal{A}^{*}(\xi),B\rangle\rvert.$
@@
    $\lvert\langle\mathcal{A}^{*}(\xi),B\rangle\rvert\leq\lVert\mathcal{A}^{*}(\xi)\rVert_{\ell_{2}}\lVert B\rVert_{*}\leq\sqrt{r^{*}}\lVert\mathcal{A}^{*}(\xi)\rVert_{\ell_{2}}.$
@@
    $\lvert\langle\mathcal{A}^{*}\mathcal{A}(E_{\lambda})-E_{\lambda},B\rangle\rvert$ $=\lvert\langle\mathcal{A}(E_{\lambda}),\mathcal{A}(B)\rangle-\langle E_{\lambda},B\rangle\rvert$
    $\leq\delta^{*}\lVert E_{\lambda}\rVert_{\mathrm{F}}\lVert B\rVert_{\mathrm{F}}$
@@
    $\lvert\langle\mathcal{A}^{*}\mathcal{A}(M_{\lambda}-\widehat{M}),B\rangle\rvert$ $\leq\lVert\mathcal{A}(M_{\lambda}-\widehat{M})\rVert\lVert\mathcal{A}(B)\rVert$
@@ 4.6 Uniqueness
(the second inequality is a standard property of RIP constants—see, e.g., [56, Proposition 6.6] for the analogous result in the sparse vector case).
Let $\widehat{M},\widehat{M}^{\prime}$ be two global optima to (2). We have shown that $\widehat{M},\widehat{M}^{\prime}$ both have rank at most $r$. Set, as before,
@@
5 Additional proofs
In this section, we state and prove two results we need on composite optimization that, though they are straightforward extensions of existing results, have not, to our knowledge, appeared as such in the literature (thus they may be of independent interest).
5.1 Convergence of constrained proximal gradient descent
Theorem 2 is an immediate consequence of the following more general result on constrained proximal gradient descent. Let $\mathcal{C}\subseteq\mathbf{R}^{d}$ be a closed set, and consider the constrained composite minimization problem
@@ 5.1 Convergence of constrained proximal gradient descent
If the minimum is not unique, we can choose one arbitrarily.
@@
In the remainder of this section, we prove Lemma 3. The proof is a modification and generalization of that of [8, Theorem 3], with many ideas (and much of our notation) taken from [57].
@@
5.2 SOCPs of factored problem are stationary for PPGD
In this section, we state and prove the below lemma, which is a straightforward extension of [9, Theorem 2.3(a)] to the case of nuclear norm regularization. A slightly weaker version of this result appeared in the blog post [58], which was itself in part inspired by an earlier version of our paper.
For twice-differentiable $f$, we denote by $\nabla^{2}f(x)[\dot{x},\dot{x}]$ the Hessian quadratic form of $f$ at $x$ applied to the direction $\dot{x}$.
Lemma 4.
Let $f\colon\mathbf{R}^{d_{1}\times d_{2}}\to\mathbf{R}$ be twice continuously differentiable. Fix a positive integer $r\leq\min\{d_{1},d_{2}\}$. For $\lambda>0$, consider the factored low-rank nuclear-norm–regularized optimization problem
$\min_{\begin{subarray}{c}X\in\mathbf{R}^{d_{1}\times r}\\ Y\in\mathbf{R}^{d_{2}\times r}\end{subarray}}~{}f(XY^{T})+\frac{\lambda}{2}(\lVert X\rVert_{\mathrm{F}}^{2}+\lVert Y\rVert_{\mathrm{F}}^{2}).$ (31)
Suppose, furthermore, that $f$ has bounded Hessian in the following sense: there is some $L>0$ such that, for all $A\in\mathbf{R}^{d_{1}\times d_{2}}$ of rank at most $r$ and $\dot{A}\in\mathbf{R}^{d_{1}\times d_{2}}$ of rank at most $2$,
$\nabla^{2}f(A)[\dot{A},\dot{A}]\leq L\lVert A\rVert_{\mathrm{F}}^{2}.$
Then, if $0<\eta\leq\frac{1}{L}$, every second-order critical point $(X,Y)$ of the factored problem (31) yields a fixed point of the rank-$r$ PPGD iteration map with stepsize $\eta$ in the sense that $M\coloneqq XY^{T}$ satisfies
$M\in\operatornamewithlimits{arg~{}min}_{\begin{subarray}{c}M^{\prime}\in\mathbf{R}^{d_{1}\times d_{2}}\\ \operatorname{rank}(M^{\prime})\leq r\end{subarray}}~{}\langle\nabla f(M),M^{\prime}\rangle+\frac{1}{2\eta}\lVert M^{\prime}-M\rVert_{\mathrm{F}}^{2}+\lambda\lVert M^{\prime}\rVert_{*}.$ (32)
The Hessian assumption could easily be relaxed to a locally Lipschitz gradient assumption like in [9], but we omit this generalization for simplicity and because we do not need it.
With this lemma, Theorem 3 follows immediately from Theorem 2. Indeed, $(2r,\delta)$-RIP implies the Hessian condition with $L=1+\delta$, and the stepsize $\eta$ used in Theorem 2 is precisely $\eta=\frac{1}{1+\delta}=\frac{1}{L}$.
Proof of Lemma 4.
Let $(X,Y)$ be a second-order critical point of (31), and set $M=XY^{T}$. By standard calculations (see, e.g., [10] or, for the part corresponding to $f$, [9]), this corresponds to the following properties. First, the gradient of the composite objective function in $(X,Y)$ is zero, that is,
$\nabla f(M)Y$ $=-\lambda X\text{ and}$
$\nabla f(M)^{T}X$ $=-\lambda Y.$
Second, the Hessian is positive semidefinite, which means, for all $\dot{X}\in\mathbf{R}^{d_{1}\times r},\dot{Y}\in\mathbf{R}^{d_{2}\times r}$,
$\langle\nabla f(M),\dot{X}\dot{Y}^{T}\rangle+\frac{1}{2}\nabla^{2}f(M)[X\dot{Y}^{T}+\dot{X}Y^{T},X\dot{Y}^{T}+\dot{X}Y^{T}]+\frac{\lambda}{2}(\lVert\dot{X}\rVert_{\mathrm{F}}^{2}+\lVert\dot{Y}\rVert_{\mathrm{F}}^{2})\geq 0.$
As $\lambda>0$, the zero-gradient conditions imply that the pair $(X,Y)$ is balanced (see, e.g., [10] for more details). Let $r_{M}=\operatorname{rank}(M)$, and let $M=U\Sigma V^{T}$ be a compact singular value decomposition of $M$ with $U^{T}U=V^{T}V=I_{r_{M}}$ and diagonal $\Sigma\in\mathbf{R}^{r_{M}\times r_{M}}$ with strictly positive diagonal entries. Then, for some $r\times r_{M}$ matrix $R$ satisfying $R^{T}R=I_{r_{M}}$, we have $X=U\Sigma^{1/2}R^{T}$ and $Y=V\Sigma^{1/2}R^{T}$; these are valid singular value decompositions of $X$ and $Y$, and we see that $X$ and $Y$ have the same singular values and right singular vectors. Furthermore, multiplying the zero-gradient conditions on the right by $R\Sigma^{-1/2}$, we obtain
$\nabla f(M)V$ $=-\lambda U\text{ and}$
$\nabla f(M)^{T}U$ $=-\lambda V.$
We can then write
$\nabla f(M)=-\lambda UV^{T}+W^{\perp},$
where $W^{\perp}$ is a matrix satisfying $W^{\perp}V=0$ and $U^{T}W^{\perp}=0$.
This is almost all we need to show (32) (cf. [9, eq. (2.4)]). The only additional optimality condition we need is
$\lVert W^{\perp}\rVert_{\ell_{2}}\leq\lambda+\frac{1}{\eta}\sigma_{r}(M).$ (33)
To show this, we apply the Hessian condition. Due to the fact that $(X,Y)$ is balanced as discussed above, there exists a unit-norm $v\in\mathbf{R}^{r}$ such that $\lVert Xv\rVert^{2}=\lVert Yv\rVert^{2}=\sigma_{r}(M)$. For arbitrary unit-norm $x\in\operatorname{range}(U)^{\perp}$, $y\in\operatorname{range}(V)^{\perp}$, taking $\dot{X}=xv^{T}$, $\dot{Y}=yv^{T}$ in the Hessian condition gives
00 0 $\leq\langle\nabla f(M),xy^{T}\rangle+\frac{1}{2}\nabla^{2}f(M)[Xvy^{T}+x(Yv)^{T},Xvy^{T}+x(Yv)^{T}]+\lambda$
$\leq\langle\nabla f(M),xy^{T}\rangle+\frac{L}{2}\lVert Xvy^{T}+x(Yv)^{T}\rVert_{\mathrm{F}}^{2}+\lambda$
$=\langle W^{\perp},xy^{T}\rangle+L\sigma_{r}(M)+\lambda.$
As this holds for any $x\in\operatorname{range}(U)^{\perp}$, $y\in\operatorname{range}(V)^{\perp}$, we must have $\lVert W^{\perp}\rVert_{\ell_{2}}\leq\lambda+L\sigma_{r}(M)$. The condition $\eta\leq\frac{1}{L}$ ensures the optimality condition (33) holds, so the result (32) follows.